# Excited $\Omega$ hyperon in charmful $\Omega_{b}$ weak decays

### v1

@@ Abstract
We investigate the sextet $b$-baryon decay processes $\Omega_{b}\to J/\psi\Omega^{(*)}$, where $\Omega^{*}$ represents the $1P$-, $1D$- and $2S$-wave excited $\Omega$ hyperons in the spectroscopy. Using the constituent quark model, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega)=2.1\times 10^{-4}$, which agrees with the previous studies to the order of magnitude. By identifying $\Omega(2012)$ as $\Omega(1^{2}P_{3/2^{-}})$, ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=2.7\times 10^{-4}$ can be similarly significant. Additionally, $\Omega(1^{2}D_{5/2^{+}})$ and $\Omega(1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}})$ states exhibit production rates of 0.6, and (0.7, 1.0), respectively, relative to their ground-state counterpart. Notably, our findings suggest that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))$ are as large as $(1.2,5.1)\times 10^{-4}$, making them accessible to experiments at LHCb.
@@ I Introduction
The $\Omega$ hyperon spectroscopy has been relatively underexplored. Nonetheless, Belle initiated a new era of discoveries with the identification of an excited hyperon, $\Omega(2012)^{-}$ Belle:2018mqs , which was later reconfirmed through the process $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Additionally, a heavier excited hyperon $\Omega(2109)^{-}$, observed via $e^{+}e^{-}\to\Omega(2109)^{-}\bar{\Omega}^{+}+c.c.$, and reported by BESIII BESIII:2024eqk . To investigate the nature of these new excited hyperons, significant theoretical attention has been devoted Xiao:2018pwe ; Liu:2019wdr ; Wang:2018hmi ; Zeng:2020och ; Aliev:2018yjo ; Aliev:2018syi ; Arifi:2022ntc ; Polyakov:2018mow ; Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh . Notably, $\Omega(2012)^{-}$ is often preferred as an exotic molecule candidate Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh rather than as a conventional $1P$-wave $\Omega(sss)$ state.
@@
Clearly, the “missing excited baryon” problem persists in $\Omega$ hyperon spectroscopy, as all waves of $\Omega$ hyperons have been theoretically predicted Oh:2007cr ; Capstick:1986ter ; Faustov:2015eba ; Loring:2001ky ; Liu:2007yi ; Chao:1980em ; Chen:2009de ; An:2013zoa ; Kalman:1982ut ; Pervin:2007wa ; An:2014lga ; Engel:2013ig ; CLQCD:2015bgi ; Carlson:2000zr ; Goity:2003ab ; Schat:2001xr ; Matagne:2006zf ; Bijker:2000gq ; Aliev:2016jnp , but only a few have been experimentally observed. These include $\Omega^{-}\equiv\Omega(1672)^{-}$ Abrams:1964tu ; Barnes:1964pd , $\Omega(2012,2109)^{-}$ Belle:2018mqs ; Belle:2021gtf ; BESIII:2024eqk , and $\Omega(2250,2380,2470)^{-}$ Biagi:1985rn ; Aston:1987bb ; Aston:1988yn . Since further clarification and exploration are essential, we propose that recent measurements of the sextet $b$-baryon decay process $\Omega_{b}^{-}\to J/\psi\Omega^{-}$ D0:2008sbw ; CDF:2009sbo ; LHCb:2013wmn ; CDF:2014mon ; LHCb:2023qxn ; Nicolini:2023stq be extended to include $\Omega_{b}^{-}\to J/\psi\Omega^{*-}$, where $\Omega^{*-}$ denotes higher-wave excited $\Omega^{-}$ hyperons.
@@
As LHCb continues to improve the statistics and precision LHCb:2023qxn ; Nicolini:2023stq , the absolute branching fractions of $\Omega_{b}^{-}$ decays may soon be measured with a similar accuracy to those of the anti-triplet $b$-baryon decays UA1:1991vse ; CDF:1992lrw ; CDF:2006eul ; CDF:1996rvy ; D0:2004quf ; D0:2007giz ; D0:2011pqa ; D0:2012hfl ; LHCb:2013hzx ; LHCb:2020iux ; LHCb:2019fim ; LHCb:2019aci ; ATLAS:2014swk ; ATLAS:2015hik ; CMS:2018wjk ; D0:2007gjs . Theoretical estimations are therefore needed, and a variety of theoretical tools are already available for such studies. These include the non-relativistic quark model Fayyazuddin:1998ap ; Mott:2011cx ; Cheng:1995fe ; Cheng:1996cs , the covariant confined quark model Gutsche:2018utw ; Gutsche:2017wag ; Gutsche:2013oea ; Gutsche:2015lea , the covariant oscillator quark model Mohanta:1998iu , the relativistic three-quark model Ivanov:1997hi ; Ivanov:1997ra , the light-front quark model Hsiao:2021mlp ; Wei:2009np ; Zhu:2018jet , the perturbative QCD approach Rui:2023fiz ; Chou:2001bn , the generalized factorization approach Hsiao:2015cda ; Fayyazuddin:2017sxq ; Hsiao:2015txa , and SU(3) flavor analysis Dery:2020lbc .
@@
In this work, we employ the constituent quark model Niu:2020gjw ; Wang:2022zja ; Niu:2021qcc ; Niu:2020aoz ; Pervin:2006ie ; Pervin:2005ve to provide our estimations. As we will demonstrate, this model can be extended to study both the $\Omega_{b}^{-}\to J/\psi\Omega^{-}$ and $\Omega_{b}^{-}\to J/\psi\Omega^{*-}$ decay processes, enabling a systematic analysis of the missing excited $\Omega^{-}$ states. This paper is organized as follows: In Sec. II, we apply the constituent quark model to two-body nonleptonic weak decays of $\Omega_{b}^{-}\to J/\psi\Omega^{(*)-}$. In Sec. III, we present the numerical results. Finally, we provide our discussions and conclusion in Sec. IV.
@@
As shown in Fig. 1, the recently observed sextet $b$-baryon decay, $\Omega_{b}^{-}\to J/\psi\Omega^{-}$, proceeds exclusively via the internal $W$-boson emission diagram Cheng:2021qpd , where $\Omega^{-}$ is strait-forwardly formed through the $\Omega_{b}^{-}\to\Omega^{-}$ transition. The single decay topology, extended to $\Omega_{b}^{-}\to J/\psi\Omega^{*-}$, also facilitates the investigation of the $1P$-, $2S$- and $1D$-wave excited $\Omega^{*-}$ states.
The quark-level effective Hamiltonian of $b\to c\bar{c}s$ weak transition induces the doubly charmful decay channels $\Omega_{b}\to J/\psi\Omega^{(*)-}$, given by Buchalla:1995vs
@@
where $G_{F}$ is the Fermi constant, $C_{1}$ and $C_{2}$ are the Wilson coefficients, and $V_{ij}$ the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements. Additionally, ${\cal O}_{1,2}$ are the current-current operators, which are written as
@@
Here, $\psi_{j_{\delta}}$ is the $j$th quark field, where $j$ can be $s$, $c$ or $b$, and $\delta=(\alpha,\beta)$ is the color index.
Figure 1: Feynman diagram for the nonleptonic weak decay $\Omega_{b}\rightarrow J/\psi\Omega^{(*)-}$.
According to Ref. Mannel:1992ti , $C_{2}\langle J/\psi\Omega^{-}|{\cal O}_{2}|\Omega_{b}^{-}\rangle$ corresponds to the internal $W$-boson emission diagram. On the other hand, $C_{1}\langle J/\psi\Omega^{-}|{\cal O}_{1}|\Omega_{b}^{-}\rangle$ is replaced by $C_{1}/N_{c}^{eff}\langle J/\psi\Omega^{-}|{\cal O}_{2}|\Omega_{b}^{-}\rangle$ in the generalized factorization, where $N_{c}^{eff}$ is an effective color number that account for the non-factorizable QCD corrections. In the large $N_{c}^{eff}$ limit ($N_{c}^{eff}\to\infty$Mannel:1992ti , the contribution from $C_{1}\mathcal{O}_{1}$ becomes negligible. Therefore, we only consider $C_{2}\mathcal{O}_{2}$ in this work.
@@
In the non-relativistic approximation, the two components are given by
$H_{W}^{PC}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}\frac{\hat{\phi}_{c}C_{2}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})\{\langle s_{3}^{\prime}|I|s_{3}\rangle$
$\langle s_{5}\bar{s}_{4}|\mbox{\boldmath$\sigma$\unboldmath}|0\rangle\left(\frac{\textbf{p}_{5}}{2m_{5}}+\frac{\textbf{p}_{4}}{2m_{4}}\right)-\Big{[}\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)\langle s_{3}^{\prime}|I|s_{3}\rangle$
$-i\langle s_{3}^{\prime}|\mbox{\boldmath$\sigma$\unboldmath}|s_{3}\rangle\times\left(\frac{\textbf{p}_{3}}{2m_{3}}-\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}\right)\Big{]}\langle s_{5}\bar{s}_{4}|\mbox{\boldmath$\sigma$\unboldmath}|0\rangle$
$-\langle s_{3}^{\prime}|I|s_{3}\rangle\Big{[}\left(\frac{\textbf{p}_{5}}{2m_{5}}+\frac{\textbf{p}_{4}}{2m_{4}}\right)\langle s_{5}\bar{s}_{4}|\mbox{\boldmath$\sigma$\unboldmath}|0\rangle-i\langle s_{5}\bar{s}_{4}|\mbox{\boldmath$\sigma$\unboldmath}|0\rangle$
×$\times\left(\frac{\textbf{p}_{4}}{2m_{4}}-\frac{\textbf{p}_{5}}{2m_{5}}\right)\Big{]}+\langle s_{3}^{\prime}|I|s_{3}\rangle\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)$
$\langle s_{5}\bar{s}_{4}|I|0\rangle\}\hat{\alpha}_{3}^{-}\,,$
$H_{W}^{PV}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}\frac{\hat{\phi}_{c}C_{2}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})\{-\langle s_{3}^{\prime}|I|s_{3}\rangle$ (4)
$\langle s_{5}\bar{s}_{4}|I|0\rangle-\langle s_{3}^{\prime}|\mbox{\boldmath$\sigma$\unboldmath}|s_{3}\rangle\langle s_{5}\bar{s}_{4}|\mbox{\boldmath$\sigma$\unboldmath}|0\rangle)\hat{\alpha}_{3}^{-}\,,$
where $\textbf{p}_{j}$ and $m_{j}$ stand for the momentum and mass of the $j$th quark, respectively, as assigned in Fig. 1, and $I$ is the dimension-two unit matrix. Additionally, $\hat{\alpha}_{3}^{-}$ is the flavor operator that transforms the quark $b$ to $s$ in the $\Omega_{b}\to\Omega^{(*)}$ transition. $s_{j}$ is the spin of the $j$rd quark. Specifically, $s^{\prime}_{3}$ and $\bar{s}_{4}$ represent the spins of the 3rd quark $c$ and the $4$th antiquark $\bar{c}$, respectively, in the $J/\psi$ production. In Eq. (II), the spin matrix elements, calculated in Niu:2020gjw ; Wang:2022zja , have been adopted. We explicitly express $\hat{\phi}_{c}$ the color operator to be
$\hat{\phi}_{c}$ $=$ $\frac{1}{3}[\hat{a}_{c}^{{\dagger}}(R)\hat{a}_{\bar{c}}^{{\dagger}}(\bar{R})+\hat{a}_{c}^{{\dagger}}(B)\hat{a}_{\bar{c}}^{{\dagger}}(\bar{B})+\hat{a}_{c}^{{\dagger}}(G)\hat{a}_{\bar{c}}^{{\dagger}}(\bar{G})]$ (5)
$[\hat{a}_{s}^{{\dagger}}(R)\hat{a}_{b}(R)+\hat{a}_{s}^{{\dagger}}(B)\hat{a}_{b}(B)+\hat{a}_{s}^{{\dagger}}(G)\hat{a}_{b}(G)]\,,$
where the $\hat{a}_{q_{1}}^{{\dagger}}(\delta)$ represents the creation operator for the quark $q_{1}$, and $\hat{a}_{q_{2}}(\delta)$ the annihilation operator for the quark $q_{2}$.
@@
$\mathcal{M}(\Omega_{b}^{-}\to J/\psi\Omega^{(*)-})=\mathcal{M}^{PC}+\mathcal{M}^{PV}\,,$ (6)
@@
$\mathcal{M}^{PC(PV)}=\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|$
$H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle\,.$ (7)

### v4

@@ Abstract
We investigate the sextet $b$-baryon decay processes $\Omega_{b}\to J/\psi\Omega^{(*)}$, where $\Omega^{*}$ represents the $1P$-, $1D$- and $2S$-wave excited $\Omega$ hyperons in the spectroscopy. Using the constituent quark model, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega)=8.8\times 10^{-4}$, which agrees with the previous studies to the order of magnitude. By identifying $\Omega(2012)$ as $\Omega(1^{2}P_{3/2^{-}})$, ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$ can be similarly significant. Additionally, $\Omega(1^{2}D_{5/2^{+}})$ and $\Omega(1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}})$ states exhibit production rates of 0.5, and (0.6, 0.8), respectively, relative to their ground-state counterpart. Notably, our findings suggest that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))$ are as large as $(4.5,20)\times 10^{-4}$, making them accessible to experiments at LHCb.
@@ I Introduction
The $\Omega$ hyperon spectroscopy has been relatively underexplored. Nonetheless, Belle initiated a new era of discoveries with the identification of an excited hyperon, $\Omega(2012)$ Belle:2018mqs , which was later reconfirmed through the process $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Additionally, a heavier excited hyperon $\Omega(2109)^{-}$, observed via $e^{+}e^{-}\to\Omega(2109)^{-}\bar{\Omega}^{+}+c.c.$, and reported by BESIII BESIII:2024eqk . To investigate the nature of these new excited hyperons, significant theoretical attention has been devoted Xiao:2018pwe ; Liu:2019wdr ; Wang:2018hmi ; Zeng:2020och ; Aliev:2018yjo ; Aliev:2018syi ; Arifi:2022ntc ; Polyakov:2018mow ; Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh . Notably, $\Omega(2012)$ is often preferred as an exotic molecule candidate Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh rather than as a conventional $1P$-wave $\Omega(sss)$ state.
@@
Clearly, the “missing excited baryon” problem persists in $\Omega$ hyperon spectroscopy, as all waves of $\Omega$ hyperons have been theoretically predicted Oh:2007cr ; Capstick:1986ter ; Faustov:2015eba ; Loring:2001ky ; Liu:2007yi ; Chao:1980em ; Chen:2009de ; An:2013zoa ; Kalman:1982ut ; Pervin:2007wa ; An:2014lga ; Engel:2013ig ; CLQCD:2015bgi ; Carlson:2000zr ; Goity:2003ab ; Schat:2001xr ; Matagne:2006zf ; Bijker:2000gq ; Aliev:2016jnp , but only a few have been experimentally observed. These include $\Omega\equiv\Omega(1672)$ Abrams:1964tu ; Barnes:1964pd , $\Omega(2012,2109)$ Belle:2018mqs ; Belle:2021gtf ; BESIII:2024eqk , and $\Omega(2250,2380,2470)$ Biagi:1985rn ; Aston:1987bb ; Aston:1988yn . Since further clarification and exploration are essential, we propose that recent measurements of the sextet $b$-baryon decay process $\Omega_{b}\to J/\psi\Omega$ D0:2008sbw ; CDF:2009sbo ; LHCb:2013wmn ; CDF:2014mon ; LHCb:2023qxn ; Nicolini:2023stq be extended to include $\Omega_{b}\to J/\psi\Omega^{*}$, where $\Omega^{*}$ denotes higher-wave excited $\Omega$ hyperons.
@@
As LHCb continues to improve the statistics and precision LHCb:2023qxn ; Nicolini:2023stq , the absolute branching fractions of $\Omega_{b}^{-}$ decays may soon be measured with a similar accuracy to those of the anti-triplet $b$-baryon decays UA1:1991vse ; CDF:1992lrw ; CDF:2006eul ; CDF:1996rvy ; D0:2004quf ; D0:2007giz ; D0:2011pqa ; D0:2012hfl ; LHCb:2013hzx ; LHCb:2020iux ; LHCb:2019fim ; LHCb:2019aci ; ATLAS:2014swk ; ATLAS:2015hik ; CMS:2018wjk ; D0:2007gjs . Theoretical estimations are therefore needed, and a variety of theoretical tools are already available for such studies. These include the non-relativistic quark model Fayyazuddin:1998ap ; Mott:2011cx ; Cheng:1995fe ; Cheng:1996cs , the covariant confined quark model Gutsche:2018utw ; Gutsche:2017wag ; Gutsche:2013oea ; Gutsche:2015lea , the covariant oscillator quark model Mohanta:1998iu , the relativistic three-quark model Ivanov:1997hi ; Ivanov:1997ra , the light-front quark model Hsiao:2021mlp ; Wei:2009np ; Zhu:2018jet ; Wang:2024mjw , the perturbative QCD approach Rui:2023fiz ; Chou:2001bn , the generalized factorization approach Hsiao:2015cda ; Fayyazuddin:2017sxq ; Hsiao:2015txa , and SU(3) flavor analysis Dery:2020lbc .
@@
In this work, we employ the constituent quark model Niu:2020gjw ; Niu:2025lgt ; Wang:2022zja ; Niu:2021qcc ; Niu:2020aoz ; Pervin:2006ie ; Pervin:2005ve to provide our estimations. As we will demonstrate, this model can be extended to study both the $\Omega_{b}\to J/\psi\Omega$ and $\Omega_{b}\to J/\psi\Omega^{*}$ decay processes, enabling a systematic analysis of the missing excited $\Omega$ states. This paper is organized as follows: In Sec. II, we apply the constituent quark model to two-body nonleptonic weak decays of $\Omega_{b}\to J/\psi\Omega^{(*)}$. In Sec. III, we present the numerical results. Finally, we provide our discussions and conclusion in Sec. IV.
@@
As shown in Fig. 1, the recently observed sextet $b$-baryon decay, $\Omega_{b}\to J/\psi\Omega$, proceeds exclusively via the internal $W$-boson emission diagram Cheng:2021qpd , where $\Omega$ is strait-forwardly formed through the $\Omega_{b}\to\Omega$ transition. The single decay topology, extended to $\Omega_{b}\to J/\psi\Omega^{*}$, also facilitates the investigation of the $1P$-, $2S$- and $1D$-wave excited $\Omega^{*}$ states.
The quark-level effective Hamiltonian of $b\to c\bar{c}s$ weak transition induces the doubly charmful decay channels $\Omega_{b}\to J/\psi\Omega^{(*)}$, given by Buchalla:1995vs
@@
where $G_{F}$ is the Fermi constant, $C_{1}$ and $C_{2}$ are the Wilson coefficients, and ($V_{bc}$, $V_{cs}$) are the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements. Additionally, ${\cal O}_{1,2}$ are the current-current operators, which are written as
@@
Here, $\psi_{j_{\delta}}$ is the $j$th quark field, where $j$ can be $s$, $c$ or $b$, and $\delta=(\alpha,\beta)$ is the color index. We define the convention for the quark and antiquark fields
$\psi_{c}(x)=\int\frac{dp}{(2\pi)^{3/2}}\left(\frac{m}{p^{0}}\right)^{1/2}\sum_{s}\left[u_{s}(p)b_{s,c}(p)e^{-ip\cdot x}+v_{s}(p)d^{\dagger}_{s,c}(p)e^{ip\cdot x}\right],$ (3)
$\bar{\psi}_{c}(x)=\int\frac{dp}{(2\pi)^{3/2}}\left(\frac{m}{p^{0}}\right)^{1/2}\sum_{s}\left[\bar{u}_{s}(p)b^{\dagger}_{s,c}(p)e^{ip\cdot x}+\bar{v}_{s}(p)d_{s,c}(p)e^{-ip\cdot x}\right].$
where $c$ is color quantum number, and
$u_{s}(p)=\sqrt{\frac{E+m}{2m}}\left(\begin{array}[]{c}\varphi_{s}\cr\frac{\mathbf{\sigma}\cdot\emph{{p}}}{E+m}\varphi_{s}\end{array}\right),v_{s}(p)=\sqrt{\frac{E+m}{2m}}\left(\begin{array}[]{c}\frac{\mathbf{\sigma}\cdot\emph{{p}}}{E+m}\chi_{s}\cr\chi_{s}\end{array}\right).$ (8)
Here, $s$ represents spin labeling, so
$\sum_{s}\varphi_{s}\varphi_{s}^{{\dagger}}=\sum_{s}\chi_{s}\chi_{s}^{{\dagger}}=1.$ (9)
The anticommutation relations of the creation and annihilation operators are given by
$\{b_{s,c}(p),b_{s^{\prime},c^{\prime}}^{\dagger}(p^{\prime})\}=\{d_{s,c}(p),d_{s^{\prime},c^{\prime}}^{\dagger}(p^{\prime})\}=\delta_{ss^{\prime}}\delta_{cc^{\prime}}\delta^{3}(p-p^{\prime}).$ (10)
The normalization of spinor is
$\bar{u}_{s}(p)u_{s^{\prime}}(p)=-\bar{v}_{s}(p)v_{s^{\prime}}(p)=\delta_{ss^{\prime}}.$ (11)
Figure 1: Feynman diagram for the nonleptonic weak decay $\Omega_{b}\rightarrow J/\psi\Omega^{(*)}$.
According to Refs. Mannel:1992ti ; Buras:1985xv ; Blok:1992na , $C_{2}\langle J/\psi\Omega|{\cal O}_{2}|\Omega_{b}\rangle$ corresponds to the internal $W$-boson emission diagram. On the other hand, $C_{1}\langle J/\psi\Omega|{\cal O}_{1}|\Omega_{b}\rangle$ is replaced by $C_{1}/N_{c}^{eff}\langle J/\psi\Omega|{\cal O}_{2}|\Omega_{b}\rangle$ in the generalized factorization, where $N_{c}^{eff}$ is an effective color number that accounts for the non-factorizable QCD corrections. In the large $N_{c}^{eff}$ limit ($N_{c}^{eff}\to\infty$Mannel:1992ti ; Buras:1985xv ; Blok:1992na , the contribution from $C_{1}\mathcal{O}_{1}$ becomes negligible. Therefore, we only consider $C_{2}\mathcal{O}_{2}$ in this work.
@@
In the non-relativistic approximation, the two components are given by Niu:2020gjw ; Niu:2025lgt ; Wang:2022zja
$H_{W}^{PC}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}C_{2}\frac{\hat{\phi}_{c}\hat{O}_{f}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})$
[…]
$-\left[\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)-i\mbox{\boldmath$\sigma$\unboldmath}_{3}\times\left(\frac{\textbf{p}_{3}}{2m_{3}}-\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}\right)\right]\cdot\mbox{\boldmath$\sigma$\unboldmath}_{4}$
[…]
$H_{W}^{PV}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}C_{2}\frac{\hat{\phi}_{c}\hat{O}^{f}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})$ (13)
[…]
Explicitly, we present
$\hat{O}_{f}$ $=$ $\hat{b}_{3^{\prime}}^{\dagger}(c)\hat{b}_{4}^{\dagger}(\bar{c})\hat{b}_{5}^{\dagger}(s)\hat{b}_{3}(b)\,,$
$\hat{\phi}_{c}$ $=$ $\delta_{c_{4}c_{3^{\prime}}}\delta_{c_{5}c_{3}}\,,$ (14)
where $\hat{b}^{({\dagger})}(q)$ annihilates (creates) the quark (q), and the $\delta_{c_{5}c_{3}}(\delta_{c_{4}c_{3^{\prime}}})$ represents that the colors of $c_{3}(c_{3^{\prime}})$ and $c_{5}(c_{4})$ should be conserved.
For the color operator to act on the color wave functions of $\Omega_{b}$, $\Omega^{(*)}$, and $J/\psi$, we define
$|\zeta_{\Omega_{b}}\rangle=\frac{1}{\sqrt{6}}(RGB-RBG+GBR-GRB+BRG-BGR)\,,$
$|\zeta_{\Omega^{(*)}}\rangle=|\zeta_{\Omega_{b}}\rangle\,,$
$|\zeta_{J/\psi}\rangle=\frac{1}{\sqrt{3}}(R\bar{R}+G\bar{G}+B\bar{B})\,.$ (15)
From Eqs. (II) and (II), we obtain
$\langle\zeta_{J/\psi}\zeta_{\Omega^{(*)}}|\hat{\phi}_{c}|\zeta_{\Omega_{b}}\rangle=\sqrt{3}\,.$ (16)
@@
$\mathcal{M}(\Omega_{b}\to J/\psi\Omega^{(*)})=\mathcal{M}^{PC}+\mathcal{M}^{PV}\,,$ (17)
@@
$\mathcal{M}^{PC(PV)}$ $=$ $\frac{1}{\sqrt{3}}\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|$ (18)
$(1+\hat{P}_{25}+\hat{P}_{15})H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle\,.$